OpenStudy (anonymous):

PLZZZ HELP!!!!!!! On a school map, the points (8, -4), (8, -10), (2, -4), and (2, -10) represent the four corners of a sand park. What are the coordinates of the center of the sand park?

OpenStudy (anonymous):

(6, -9) (5, -7) (4, -2) (8, -5)

OpenStudy (anonymous):

the center is the midpoint of the line joining points (2, -4) and (8 ,-10)

OpenStudy (anonymous):

Im still confused?

OpenStudy (anonymous):

the figure is a rectangle and midpoint is middle of one of the diagonals

OpenStudy (anonymous):

so what would be the answer?

OpenStudy (anonymous):

midpoint = ( 8+2) (-4-10) --- , ------ 2 2 - simplify and youve got it

OpenStudy (anonymous):

how do you simplify? im sorry i got stuck into doing my older brothers homework =/

OpenStudy (anonymous):

ooohh (10/2 . -14/2) = (5, -7)

OpenStudy (anonymous):

thank you i really appreciate it! if you dont mind could you help me with a couple more?

OpenStudy (anonymous):

plz press best answer

OpenStudy (anonymous):

and repost further questions

OpenStudy (anonymous):

A scale drawing of a storage box on a coordinate grid shows a square with the points (-3, 5) and (1, 7) as the locations of two adjacent vertices. Which pair of coordinate points could represent the other two vertices?

OpenStudy (anonymous):

(0, 1) and (3, 3) (-5, 9) and (1, 11) (-5, 9) and (-1, 11) (0, 1) and (4, 3)